# The impact of assembly history on the X-ray detectability of halos

### v1

@@ 4 Discussion
Different studies have investigated the dependence of halo properties such as their formation time, accretion rate, spin, shape, velocity dispersion, concentration, and anisotropy (Gao et al. 2005; Gao & White 2007; Faltenbacher & White 2009; Borzyszkowski et al. 2017; Ganeshaiah Veena et al. 2018; Paranjape et al. 2018) with the environment. While this dependence is not unexpected (Sheth & Tormen 2004; Avila-Reese et al. 2005; Gao et al. 2005; Gao & White 2007; Weinmann et al. 2006), the specific role of the non-linear cosmic web in this matter is still highly debated and not completely understood. However, our result connecting the preferential location of X-ray bright halos in denser environments aligns with previous findings that highlight the role of large scale structures in modulating halo gas content in different structures (Romano-Diaz et al. 2017; Donnan et al. 2022; Gouin et al. 2022, 2023; Davies 2023; Ma et al. 2025; Hasan et al. 2024; Zarattini et al. 2024).
@@ 4.2 A look out into the future
However, assigning galaxies to groups can introduce significant biases in velocity dispersion measurements (e.g. Old et al. 2015), potentially affecting mass estimates. Optical methods are also highly susceptible to contamination when reconstructing total group and cluster masses. Marini et al. (2025a) highlighted the scatter in halo mass estimates when using velocity dispersion-based definitions. It is in our plan to revisit the stacking procedure outlined in Popesso et al. (2024a) taking these effects into account.
@@
To address these challenges, future work will refine the stacking procedure and DM concentration reconstruction to mitigate such biases. Our primary goal is to assess whether environment-driven accretion and halo assembly bias significantly shape the baryonic content of galaxy groups and clusters, beyond the effects of AGN feedback. When these effects are properly accounted for, stacking can be a powerful tool, even enabling the recovery of spectral-reconstructed properties such as halo temperature (Toptun et al. submit). A step further can be taken by reconstructing filaments and voids in the large scale environment and investigating these properties as a function of the distance from the filaments with tools such as DisPerSE (Sousbie 2011; Sousbie et al. 2011) and the Monte Carlo Physarum Machine (MCPM; Hasan et al. 2024), to name a few. However, this goes beyond the scope of this paper and we leave this application for future work.
@@
Additionally, since X-ray emission traces hot gas and is highly sensitive to density, it primarily reflects the most concentrated regions of the IGrM. However, the SZ effect (Zeldovich & Sunyaev 1969; Sunyaev & Zeldovich 1972), which depends on the integrated pressure, can provide complementary insights – especially in cases where AGN feedback or mergers redistribute the gas without fully expelling it. This means that while X-ray emission may drop due to gas becoming more diffuse over time, the SZ signal could persist, making it valuable for capturing the full picture of a halo’s baryon content and thermodynamic history (see Mroczkowski et al. 2019, for a review).
Hydrodynamical cosmological simulations suggest that the integrated SZ signal $Y_{\mathrm{SZ}}$ is one of the most reliable proxies for total cluster mass, exhibiting a scatter of approximately 10% (e.g. Motl et al. 2005). This has motivated the development of low-scatter X-ray mass proxies, such as the core-excised quantity $Y_{X}\equiv M_{\mathrm{gas}}T_{X}$, introduced by Kravtsov et al. (2006).
Given this, jointly analysing X-ray and SZ signals offers a more complete understanding of how feedback and accretion shape the gaseous haloes of galaxy groups and clusters. Future analysis will explore the impact of the SZ effect on recovering the baryonic content at the mass scales.
@@
Figure 10: Schematic representation of the assembly history of the bright (top) and faint (bottom) galaxy groups in simulation. The colour scheme (shown as a green-purple gradient) represents the hot gas fraction in the halo within the virial radius, while the sizes of the circles are proportional to the halo mass. At their core, groups host SMBHs (black circles) triggered by AGN episodes at different intensities (as depicted by the coloured arrows), depending on the surrounding gas content. The BH sizes are exaggerated to highlight the difference in the $M_{BH}/M_{\mathrm{halo}}$ ratio between the bright and faint samples. The colour and size of the arrow hint at the strength of the AGN activity: redder and larger arrows correspond to a stronger effect. The bright sample is characterised by a fast mass accretion at the late stages of their evolution given that they are located in dense environments, while the faint one comprises halos evolving relatively isolated. The dotted lines illustrate the typical self-similar evolution expected for a DM halo with an NFW profile, fitting the evolution of the faint sample.
@@ 5 Conclusions
In this work, we investigated the key physical processes driving biases in X-ray-selected samples of galaxy groups within the Magneticum simulations. We focused on the role of AGN feedback and halo assembly history in shaping their gas content, which in turn affects their X-ray properties. Our study is motivated by the vast amount of data eROSITA currently provides at these mass scales (Merloni et al. 2012).
@@
  • Extended luminous groups have high surface brightness without a strong central peak but still retain more hot gas than faint systems.
@@
Our analysis reveals that the AGN activity primarily depletes haloes of baryons, particularly at early times ($z\sim 2-3$), inducing lower X-ray luminosities, but late-time accretion events, such as mergers at $z<1$, can replenish the gas content, increasing the total gas fraction (see Fig. 5). This newly acquired gas fuels the central BH, leading to further AGN feedback that raises the plasma temperature and enhances X-ray emissivity, though without immediate baryon depletion (see the schematic representation in Fig 10). Whether a halo remains gas-rich or depleted depends on its merger history: halos without late-time mergers will experience continued baryon loss due to AGN feedback, leading to declining X-ray emissivity, while those undergoing gas-rich mergers can temporarily boost their X-ray brightness. Over sufficiently long timescales, even today’s bright systems may eventually be depleted by AGN activity, aligning with the faint population. This highlights the need to consider both AGN feedback and merger history (associated with the local environment, as illustrated in Fig. 7) in explaining the presence or absence of hot gas in halos, rather than attributing it solely to AGN feedback.
@@
We confirm that gas-poor halos tend to host more massive and older stellar populations (illustrated in Fig. 9), consistent with early-formed systems having a longer time for accretion and star formation. Similarly, we observe a strong anti-correlation between the hot gas fraction and the system’s fossilness, suggesting that groups with a more dominant central galaxy have undergone significant early gas depletion. Gas-poor systems exhibit enhanced metal content due to longer stellar evolution and AGN-driven redistribution of metals. AGN activity also affects the temperature, with continuous thermal feedback at late times. Additionally, we find that gas-rich halos tend to reside in denser environments, as indicated by a higher number of satellite galaxies within $R_{500}$. Finally, our results suggest that the star formation activity of central galaxies differs significantly between X-ray bright and faint systems, as a natural consequence of the gas depletion in X-ray faint groups.

### v2

@@ 4 Discussion
Different studies have investigated the dependence of halo properties such as their formation time, accretion rate, spin, shape, velocity dispersion, concentration, and anisotropy (Gao et al. 2005; Gao & White 2007; Faltenbacher & White 2009; Borzyszkowski et al. 2017; Ganeshaiah Veena et al. 2018; Paranjape et al. 2018) on the environment. While this dependence is not unexpected (Sheth & Tormen 2004; Avila-Reese et al. 2005; Gao et al. 2005; Gao & White 2007; Weinmann et al. 2006), the specific role of the non-linear cosmic web in this matter is still highly debated and not completely understood. However, our result connecting the preferential location of X-ray bright halos in denser environments aligns with previous findings that highlight the role of large-scale structures in modulating halo gas content in different structures (Romano-Diaz et al. 2017; Donnan et al. 2022; Gouin et al. 2022, 2023; Davies 2023; Ma et al. 2025; Hasan et al. 2024; Zarattini et al. 2024).
@@ 4.2 A look out into the future
However, assigning galaxies to groups can introduce significant biases in velocity dispersion measurements (e.g. Old et al. 2015), potentially affecting mass estimates. Optical methods are also highly susceptible to contamination when reconstructing total group and cluster masses. Marini et al. (2025) highlighted the scatter in halo mass estimates when using velocity dispersion-based definitions. It is in our plan to revisit the stacking procedure outlined in Popesso et al. (2024a), taking these effects into account.
@@
To address these challenges, future work will refine the stacking procedure and DM concentration reconstruction to mitigate such biases. Our primary goal is to assess whether environment-driven accretion and halo assembly bias significantly shape the baryonic content of galaxy groups and clusters, beyond the effects of AGN feedback. When these effects are properly accounted for, stacking can be a powerful tool, even enabling the recovery of spectral-reconstructed properties such as halo temperature (Toptun et al. 2025, submitted). A step further can be taken by reconstructing filaments and voids in the large scale environment and investigating these properties as a function of the distance from the filaments with tools such as DisPerSE (Sousbie 2011; Sousbie et al. 2011) and the Monte Carlo Physarum Machine (MCPM; Hasan et al. 2024), to name a few. However, this goes beyond the scope of this paper, and we leave this application for future work.
@@
Additionally, since X-ray emission traces hot gas and is highly sensitive to density, it primarily reflects the most concentrated regions of the IGrM. However, the SZ effect (Zeldovich & Sunyaev 1969; Sunyaev & Zeldovich 1972), which depends on the integrated pressure, can provide complementary insights – especially in cases where AGN feedback or mergers redistribute the gas without fully expelling it. This means that while X-ray emission may drop due to gas becoming more diffuse over time, the SZ signal could persist, making it valuable for capturing the full picture of a halo’s baryon content and thermodynamic history (see Mroczkowski et al. 2019, for a review). Hydrodynamical cosmological simulations suggest that the integrated SZ signal $Y_{\mathrm{SZ}}$ is one of the most reliable proxies for total cluster mass, exhibiting a scatter of approximately 10% (e.g. Motl et al. 2005). This has motivated the development of low-scatter X-ray mass proxies, such as the core-excised quantity $Y_{X}\equiv M_{\mathrm{gas}}T_{X}$, introduced by Kravtsov et al. (2006). Given this, jointly analysing X-ray and SZ signals offers a more complete understanding of how feedback and accretion shape the gaseous haloes of galaxy groups and clusters. Future analysis will explore the impact of the SZ effect on recovering the baryonic content at the mass scales.
@@
Figure 10: Schematic representation of the assembly history of the bright (top) and faint (bottom) galaxy groups in the simulation. The colour scheme (shown as a green-purple gradient) represents the hot gas fraction in the halo within the virial radius, while the sizes of the circles are proportional to the halo mass. At their core, groups host SMBHs (black circles) triggered by AGN episodes at different intensities (as depicted by the coloured arrows), depending on the surrounding gas content. The BH sizes are exaggerated to highlight the difference in the $M_{BH}/M_{\mathrm{halo}}$ ratio between the bright and faint samples. The colour and size of the arrow hint at the strength of the AGN activity: redder and larger arrows correspond to a stronger effect. The bright sample is characterised by a fast mass accretion at the late stages of their evolution, given that they are located in dense environments, while the faint one comprises halos evolving relatively isolated in under-dense regions. The dotted lines illustrate the typical self-similar evolution expected for a DM halo with an NFW profile, fitting the evolution of the faint sample.
@@ 5 Conclusions
In this work, we investigated the key physical processes driving biases in X-ray-selected samples of galaxy groups within the Magneticum simulations. We focused on the role of environment, AGN feedback, and halo assembly history in shaping their gas content, which in turn affects their X-ray properties. Our study is motivated by the vast amount of data eROSITA currently provides at these mass scales (Merloni et al. 2012).
@@
  • Extended luminous groups have high surface brightness without a strong central peak, but retain more hot gas than faint systems.
@@
Our analysis reveals that the AGN activity primarily depletes haloes of baryons, particularly at early times ($z\sim 2-3$), inducing lower X-ray luminosities, but late-time accretion events, such as mergers at $z<1$, can replenish the gas content, increasing the total gas fraction (see Fig. 5). This newly acquired gas fuels the central BH, leading to further AGN feedback that raises the plasma temperature and enhances X-ray emissivity, though without immediate baryon depletion (see the schematic representation in Fig 10). Whether a halo remains gas-rich or depleted depends on its merger history: halos without late-time mergers experience continued baryon loss due to AGN feedback, leading to declining X-ray emissivity, while those undergoing gas-rich mergers can temporarily boost their X-ray brightness. Over sufficiently long timescales, even today’s bright systems may eventually be depleted by AGN activity, aligning with the faint population. This highlights the need to consider both AGN feedback and merger history (associated with the local environment, as illustrated in Fig. 7) in explaining the presence or absence of hot gas in halos, rather than attributing it solely to AGN feedback.
@@
We confirm that gas-poor halos tend to host more massive and older stellar populations (illustrated in Fig. 9), consistent with early-formed systems having a longer time for accretion and star formation. Similarly, we observe a strong anti-correlation between the hot gas fraction and the system’s fossilness, suggesting that groups with a more dominant central galaxy have undergone significant early gas depletion. Gas-poor systems exhibit enhanced metal content due to longer stellar evolution and AGN-driven redistribution of metals. AGN activity also affects the temperature, with continuous thermal feedback at late times. Additionally, we find that gas-poor halos tend to reside in less dense environments, as indicated by a lower number of satellite galaxies within $R_{500}$. Finally, our results suggest that the star formation activity of central galaxies differs significantly between X-ray bright and faint systems, as a natural consequence of the gas depletion in X-ray faint groups.